### v1

# The Non Parametric Reconstruction of Binary Black Hole Mass Evolution from GWTC-4.0 Gravitational Wave Catalog

Abstract
The distribution of binary black hole (BBH) masses and its evolution with redshift provide key insights into the different formation channels of the compact objects and their evolution with cosmic time and stellar properties such stellar metallicity and star formation rate history. We present a non parametric, model-independent joint reconstruction of the redshift evolution of BBH mass distribution from gravitational wave (GW) catalog GWTC-4.0 from the fourth observation of LIGO-Virgo-KAGRA (LVK). This method simultaneously searches for the signature of any linear and quadratic redshift evolution with respect to the low redshift in a Bayesian framework taking into accounting the detector selection effects. We find tentative evidence for a linear redshift-dependent evolution of the mass distribution, consistent over a mass range ($m\gtrsim 50\,M_{\odot}$). While lower mass systems shows no signature of evolution. The quadratic term remains consistent with zero, indicating that a simple linear dependence adequately describes the population up to redshift $z\sim 1$. In future with more GW sources, this technique can shed light into the true nature of the redshift dependence and possibility to uncover subtle evolutionary features in BBH populations and to probe the cosmic history of black hole formation.
1 Introduction
The detection of gravitational waves (GWs) has revolutionized astrophysics by providing a direct means to observe and characterize binary black hole populations across vast cosmic distances. These unprecedented observations have transformed our understanding of how stellar mass black holes form, evolve, and merge over the history of the universe (Abbott et al., 2019, 2016a; Bouffanais et al., 2021; Franciolini et al., 2022; Cheng et al., 2023; Antonelli et al., 2023; Wong et al., 2023; Tiwari & Fairhurst, 2021; Tiwari, 2022, 2025; Kimball et al., 2020, 2021; Afroz & Mukherjee, 2025a, b, c). With each new detection, we gain invaluable information about the masses, spins, and rates of these elusive objects, enabling tests of astrophysical theory in regimes previously inaccessible (Abbott et al., 2021, 2023; LIGO-VIRGO-KAGRA Collaboration, 2025a).
A fundamental question arising from these advances is whether the underlying distribution of binary black hole (BBH) masses, called as the BBH mass function changes over cosmic time. Investigating this evolution is crucial because it encodes rich information about the life cycles of massive stars, the influence of their environments, and the physical processes shaping their end states (Mukherjee, 2022; Bailes et al., 2021; Arimoto et al., 2021; Mapelli, 2021; Karathanasis et al., 2023; Iorio et al., 2022; Rinaldi et al., 2025; Barrett et al., 2018; Dominik et al., 2012; Bailyn et al., 1998; Ye & Fishbach, 2024; Schiebelbein-Zwack & Fishbach, 2024; Rinaldi et al., 2024; Gennari et al., 2025; Lalleman et al., 2025; Sicilia et al., 2022).
Factors such as the chemical composition of progenitor stars, commonly referred to as metallicity, vary significantly across epochs and profoundly impact stellar winds and mass loss rates. Lower metallicities prevalent in the early universe likely allowed stars to retain more mass, potentially leading to the formation of heavier black holes (Karathanasis et al., 2023; Farrell et al., 2021; Belczynski et al., 2019; Spera & Mapelli, 2017; Farmer et al., 2020; Hendriks et al., 2023; Farmer et al., 2019; Leung et al., 2019; Tong et al., 2025; Antonini et al., 2025b; Costa et al., 2022; Fishbach & Holz, 2017; Talbot & Thrane, 2018). Moreover, the universe’s first generation of stars, the Population III stars, constituted metal free environments that could have produced exceptionally massive black holes distinct from those observed today (Santoliquido et al., 2023; Costa et al., 2023; Giacobbo et al., 2018; Dvorkin et al., 2016).
Understanding how these astrophysical processes imprint themselves on the black hole mass function as a function of redshift holds the promise of uncovering the origins and evolution of compact object populations, offering insights into star formation history, galaxy evolution, and the end stages of stellar lives. Yet, measuring the redshift dependent mass distribution presents formidable challenges (Spera & Mapelli, 2017; Belczynski et al., 2016; Tanikawa, 2024). Previous studies have explored various approaches to characterize the BBH mass distribution and its potential evolution, ranging from parametric models with specific functional forms to more flexible non-parametric methods (Wysocki et al., 2019; Mandel et al., 2019; Tiwari & Fairhurst, 2021; Talbot et al., 2019; Wong et al., 2020; Callister & Farr, 2024; Riley & Mandel, 2023; Farah et al., 2025; Talbot et al., 2025; Ray et al., 2023; Magaña Hernandez & Palmese, 2025; Antonini et al., 2025a; Edelman et al., 2023; Heinzel et al., 2025b, 2023, a; Alvarez-Lopez et al., 2025).
Figure 1: Schematic illustration of the impact of selection effects and intrinsic mass evolution on the observable black hole mass distribution. The red regions denote parameter space inaccessible to detectors due to selection bias. The green area represents the detectable mass range under scenarios with no intrinsic evolution, while the yellow region highlights where signatures of positive mass evolution would appear. For negative evolution, part of the green region below the blue line and above the lower red boundary remains observable. Arrows and shaded bands indicate the characteristic observational trends: constant distribution for no evolution, downward shifts for negative evolution, and upward shifts for positive evolution, emphasizing how evolutionary effects manifest relative to the standard detection boundaries.
The two major open questions regarding the redshift evolution of the BBH mass distribution are: (i) Does a single, global BBH mass distribution adequately describe the population across all redshifts, or is such a description valid only within restricted redshift ranges? (ii) If not, does the BBH mass distribution itself evolve with redshift, and if so, what is the nature of that evolution? The challenge in answering these questions lies with our ignorance in modelling either of these from first principle due to large astrophysical uncertainties. Along with the theoretical uncertainties, selection effects also plays a major role.
In order to infer the true physical effect from the data, one needs to take into account the redshift and mass dependent selection effects. Gravitational wave detectors have sensitivity biases favoring the detection of more massive and nearby sources, creating a mass and redshift dependent selection function effect, which can be modeled given the detectors sensitivity (Essick, 2023; Essick & Holz, 2024; Messenger & Veitch, 2013; Gerosa & Bellotti, 2024).
Figure 1 presents a schematic illustration of the black hole mass distribution as a function of redshift, highlighting the role of GW selection effects and evolutionary signatures. The red-shaded regions indicate areas excluded from detection due to sensitivity limitations, while the yellow region represents the detectable parameter space. In the absence of intrinsic evolution, the observed black hole masses should follow a band bounded by the selection threshold and a constant upper mass limit (dashed blue line). However, the presence of mass evolution introduces systematic shifts, positive evolution leads to an increasing upper mass boundary with redshift (dashed black line), while negative evolution reduces it (dashed blue line). The figure thus encapsulates how black hole mass distributions that evolve with cosmic time can be disentangled from selection bias, providing direct observational probes of black hole formation and stellar evolution across the universe.
In this work, we propose a novel, fully non parametric framework for inferring the evolution of the stellar mass black hole mass function using GW observations. Leveraging a Taylor series expansion of the joint mass redshift distribution, this approach makes minimal assumptions about the functional form of the mass function or its evolution. Instead, it allows the data itself to reveal evolutionary patterns, capturing both linear and higher order changes in the population over redshift. This model-independent strategy offers greater flexibility and interpretability, with each Taylor coefficient corresponding to physically meaningful rates and accelerations of population change. Such a framework paves the way for revealing subtle and complex evolutionary signatures that could otherwise remain hidden. Applying this framework to the GWTC-3 (Abbott et al., 2023) and GWTC-4.0 catalog (LIGO-VIRGO-KAGRA Collaboration, 2025b), the most comprehensive collection of BBH mergers to date, we find tentative evidence for mass dependent evolution in the BBH mass function. High mass systems ($40$-$60\,M_{\odot}$) show positive tentative redshift evolution, while lighter black holes exhibit little or no statistically significant evolution. These results are consistent with metallicity driven formation channels and demonstrate the potential of model independent methods for probing black hole populations with future GW observations.
2 Non Parametric Bayesian Inference of the BBH Mass Distribution with Redshift
The aim of this work is to test whether the distribution of BBH component masses evolves with redshift, a question with direct implications for stellar evolution, metallicity, and the astrophysical environments of BBH formation. Gravitational wave observations provide unique access to this distribution, but detector sensitivity introduces strong selection effects: massive systems are detectable at greater distances than lighter ones, biasing the observed sample. Correcting for these biases is essential for reconstructing the intrinsic population. We demonstrate in these sections our Bayesian method in reconstruction of BBH mass distribution from data in a model-independent way.
2.1 Non parametric Mass Redshift Distribution
Rather than relying on parametric models with fixed functional forms, which risk obscuring unanticipated features, we adopt a non parametric, model free framework based on a Taylor expansion of the mass redshift distribution. Instead of prescribing exact shapes, this method expresses $p(m,z)$ as a smooth series expansion in redshift around a reference redshift $z_{\rm ref}$:
$p(m,z)=\sum_{n=0}^{N}\frac{1}{n!}\left.\frac{\partial^{n}p(m,z)}{\partial z^{n}}\right|_{z=z_{\rm ref}}(z-z_{\rm ref})^{n},$ (1)
where the truncation order $N$ dictates the level of evolutionary behavior captured, and we choose $z_{\rm ref}=0.25$ to center the expansion of the mass distribution. In this work, we restrict our analysis up to second order term in the Taylor series expansion, justified by the fact that the redshift range of current GW observations extends up to $z\sim 1.0$, along with only a few detected events at high redshift. The resulting approximation is:
$p(m,z)\approx p_{0}(m)+p_{1}(m)(z-z_{\rm ref})+p_{2}(m)(z-z_{\rm ref})^{2},$ (2)
with the coefficient functions defined as
$p_{0}(m)\equiv p(m,z_{\rm ref}),\\ p_{1}(m)\equiv\left.\frac{\partial p}{\partial z}\right|_{z_{\rm ref}},\\ p_{2}(m)\equiv\frac{1}{2}\left.\frac{\partial^{2}p}{\partial z^{2}}\right|_{z_{\rm ref}}.$ (3)
Here, $p_{0}(m)$ encodes the mass distribution at the reference redshift, while $p_{1}(m)$ and $p_{2}(m)$ capture the linear and quadratic evolutionary trends respectively. This construction avoids restrictive parametric assumptions, permits non-monotonic evolution, and provides coefficients with direct physical interpretations.
The recovered evolution coefficients encode distinct aspects of mass function evolution. The linear evolution coefficient $p_{1}(m)$ indicates whether black holes of mass $m$ were more common (positive values) or less common (negative values) at higher redshifts. The magnitude $|p_{1}(m)|$ quantifies the strength of evolution, with $|p_{1}(m)|\sim 1$ corresponding to order-unity changes over $\Delta z\sim 1$. The quadratic evolution coefficient $p_{2}(m)$ captures curvature in evolutionary trends. Positive values of $p_{2}(m)$ indicate accelerating evolution, while negative values suggest deceleration. This term allows for non-monotonic behavior and can reveal whether evolutionary processes are intensifying or saturating with redshift.
2.2 Data Processing and Selection Function
[…]
$d_{L}(z)=(1+z)\,\frac{c}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{m}(1+z^{\prime})^{3}+(1-\Omega_{m})}},$ (4)
here $\Omega_{m}$ is the matter density and $H_{0}$ is the Hubble constant. We convert each distance sample $d_{L,s}^{(i)}$ to a redshift $z_{s}^{(i)}$ using this relation, mapping the posteriors into the $(m,z)$ plane. We work consistently in the source frame mass distribution for easily understanding the signature of any redshift evolution in the source frame mass distribution.
To calibrate the selection function, we use injection campaigns in which artificial signals of known parameters are added to the detector noise and recovered with the search pipelines. For each injection, we record its source frame masses, luminosity distance, and network signal to noise ratio, classifying it as detected if $\mathrm{SNR}>\rho_{\mathrm{thresh}}$ with $\rho_{\mathrm{thresh}}=10$. We note that while GWTC detections are defined using a false alarm rate threshold, in the injection sets we adopt the SNR cut as the criterion for detectability.
The selection function $S(m,z)$ represents the probability that a binary with component mass $m$ at redshift $z$ will be detected by the GW network. Formally, this is the marginalization over extrinsic parameters $\Omega$ (sky location, orientation, etc.):
$S(m,z)=\int P(\text{det}|m,z,\Omega)\,p(\Omega)\,d\Omega\approx\frac{N_{\mathrm{det}}(m,z)}{N_{\mathrm{inj}}(m,z)},$ (5)
where the integral is estimated empirically from injection campaigns. The LIGO(Abbott et al., 2016b) Virgo(Acernese et al., 2015) KAGRA(Akutsu et al., 2021) (LVK) Collaboration performs extensive injection campaigns by adding simulated GW signals with known parameters to detector noise, then processing these through the same search pipelines used for actual detections. These injections span the parameter space of compact binary coalescences, including various mass ratios, spins, and orientations, providing a comprehensive mapping of detector sensitivity across the $(m,z)$ plane (Abbott et al., 2023; Essick & Holz, 2024; Talbot & Thrane, 2018; Gerosa & Bellotti, 2024; Essick et al., 2025)  [^1] The injection sets and recovered selection functions for GWTC-3 and GWTC-4.0 are publicly available at https://zenodo.org/record/5636816 and https://zenodo.org/records/16740128 respectively.. In practice, we discretize the $(m,z)$ plane into a two dimensional grid of bins $\{(m_{k},z_{j})\}$ with $k=1,\ldots,N_{m}$ and $j=1,\ldots,N_{z}$, computing $S(m_{k},z_{j})=N_{\mathrm{det}}(m_{k},z_{j})/N_{\mathrm{inj}}(m_{k},z_{j})$ where $N_{\mathrm{inj}}$ and $N_{\mathrm{det}}$ are the number of injected and detected signals in each bin. We impose a floor $S(m_{k},z_{j})\geq 0.01$ to avoid numerical instabilities in regions with sparse injections.
2.3 Expected Number of Detections
The key insight is that the observed number of detections in each bin depends on three factors: the intrinsic merger rate density, the mass distribution, and the selection effects. The expected number of detections in bin $(m_{k},z_{j})$ is:
$N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})$ $=T_{\rm obs}\int_{m_{k}}^{m_{k}+\Delta m_{k}}\int_{z_{j}}^{z_{j}+\Delta z_{j}}\mathcal{R}(z)\frac{p(m|z,\boldsymbol{\theta})}{1+z}$
$\times\frac{dV_{c}}{dz}S(m,z)\,dm\,dz,$ (6)
where $T_{\rm obs}$ is the total observation time, $\mathcal{R}(z)$ is the merger rate density per unit comoving volume at redshift $z$, $p(m|z,\boldsymbol{\theta})$ is the normalized mass distribution at redshift $z$, $(1+z)^{1}$ accounts for cosmological time dilation, and $dV_{c}/dz=4\pi cd_{L}^{2}(z)/[(1+z)^{2}H(z)]$ is the differential comoving volume per unit redshift.
For computational efficiency, we factorize this expression. Assuming the merger rate evolution can be absorbed into our Taylor expansion framework, we write:
$N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})$ $=N_{\mathrm{tot}}\cdot\tilde{p}(m_{k},z_{j}|\boldsymbol{\theta})\cdot S(m_{k},z_{j})$
$\times\frac{dV_{c}}{dz}\bigg|_{z_{j}}\cdot\frac{1}{1+z_{j}}\cdot\Delta m_{k}\Delta z_{j},$ (7)
where $N_{\mathrm{tot}}=\sum_{k,j}N_{\mathrm{obs}}(m_{k},z_{j})$ is the total number of observed events, and $\tilde{p}(m_{k},z_{j}|\boldsymbol{\theta})$ is our Taylor-expanded mass-redshift distribution. The factor $dV_{c}/dz|_{z_{j}}$ provides the comoving volume element at each redshift bin, naturally weighting the contribution of each bin by the available survey volume. The factor $(1+z_{j})^{-1}$ accounts for the cosmological time dilation.
2.4 Likelihood and Inference
The population parameters are $\boldsymbol{\theta}=\{p_{0}(m_{k}),p_{1}(m_{k}),p_{2}(m_{k})\}_{k=1}^{N_{m}}$. We require the intrinsic model to remain positive and properly normalized across all bins.
Assuming Poisson statistics for the observed counts in each bin, the likelihood is:
$\mathcal{L}(\boldsymbol{\theta})=\prod_{k,j}\frac{N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})^{N_{\mathrm{obs}}(m_{k},z_{j})}e^{N_{\mathrm{exp}}(m_{k},z_{j}|\boldsymbol{\theta})}}{N_{\mathrm{obs}}(m_{k},z_{j})!}.$ (8)
This Poisson likelihood models the observed counts in each bin denoted by the index k, and j, where $N_{\mathrm{exp}}(m_{k},z_{j})$ represents the expected number of detections given the intrinsic distribution and selection effects, while $N_{\mathrm{obs}}(m_{k},z_{j})$ are the actual observed counts. The Poisson distribution naturally accounts for the discrete nature of GW detections and the statistical fluctuations in sparse bins, making it particularly suitable for the relatively small sample sizes typical of current GW catalogs. To account for parameter estimation uncertainties, we use the full posterior sample of every event, ensuring full posterior support is explored across the ensemble. Events spanning multiple bins naturally contribute to coefficient uncertainties through this sampling procedure.
We adopt priors that ensure physical behavior while remaining minimally informative. We use uniform priors $p_{1}(m_{k})\sim\text{Uniform}(-4,4)$ and $p_{2}(m_{k})\sim\text{Uniform}(-4,4)$ that encompass physically reasonable evolutionary timescales. We sample from this posterior using ensemble Monte Carlo methods.We sample from the posterior using the emcee ensemble sampler (Foreman-Mackey et al., 2013). This framework thus provides a model independent probe of BBH mass function evolution, with coefficients that have direct astrophysical interpretations.
3 Observed BBH Population and Evidence for Mass Distance Trends
Figure 2: Exploratory visualization of observed BBH component mass and distance distributions from the GWTC-3 and GWTC-4.0 catalogs. The top left panel shows a scatter plot of median posterior component masses versus median luminosity distance for GWTC-3, categorized as low (20-40 M, blue), medium (40-60 M, purple), and high (>>60 M, orange) mass ranges, highlighting the distribution of detected systems. The top right panel presents violin plots for GWTC-3, with the left violins showing the distributions of primary component masses and the right violins showing distributions for secondary masses, grouped in three luminosity distance bins (0-2, 2-4.5, and 4.5-8 Gpc), illustrating how both mass components vary with distance. The bottom left panel repeats the mass versus luminosity distance scatter plot for GWTC-4.0, revealing analogous mass-dependent distance patterns. The bottom right panel displays violin plots for GWTC-4.0, again distinguishing primary (left) and secondary (right) mass distributions within each distance bin, revealing the changing spread and range of observed masses at increasing distances. Collectively, these panels illustrate the evolving mass-distance relationship across observing runs and motivate the application of hierarchical population analyses that rigorously account for uncertainties and selection biases. Importantly, low-mass BBHs are not detectable at higher redshifts primarily due to selection effects, while the absence of high-mass BBHs at lower distances cannot be explained by selection biases alone, suggesting intrinsic astrophysical factors influence their local detectability.
[…]
Figure 2 presents event-level summary statistics illustrating qualitative trends in the detected samples from the GWTC-3 and GWTC-4.0 catalogs. The top row shows results for GWTC-3, with the left panel displaying median component mass versus median luminosity distance for each posterior, grouped into three mass categories (low: 20-40 M, medium: 40-60 M, high: >>60 M). The right panel shows violin plots of primary (left) and secondary (right) component mass distributions within each luminosity distance bin (0-2, 2-4.5, and 4.5-8 Gpc), demonstrating how both mass components vary with distance. The bottom row provides the corresponding GWTC-4.0 results, with the left panel depicting the mass versus distance scatter plot and the right panel presenting violin plots for primary (left) and secondary (right) mass within the same luminosity distance bins. Together, these visual summaries reveal a stronger correlation between component mass and distance in GWTC-4.0 compared to GWTC-3, consistent with GWTC-4.0’s broader redshift coverage. These trends confirm that higher-mass BBHs tend to be detected at greater distances (impacted by selection effects) and exhibit broader mass distributions across distance ranges.
Such patterns can arise from two non exclusive effects. First, heavier binaries generate stronger GW signals and are therefore detectable at larger distances; this selection bias alone can produce an apparent mass redshift correlation even if the intrinsic population is non evolving. Second, there may be genuine astrophysical evolution of the BBH mass function for example, metallicity driven differences in stellar winds and mass loss, that enhances the formation of massive remnants at earlier cosmic times. Disentangling these two effects requires quantitative correction for detector sensitivity and measurement uncertainty.
Figure 3: Posterior constraints on the redshift evolution of the BBH mass distribution, shown as a function of component mass. Left: Linear coefficient $p_{1}(m)$, measuring the first-order rate of change of the mass distribution with redshift. Right: Quadratic coefficient $p_{2}(m)$, capturing any curvature in the redshift dependence. Error bars denote $1\sigma$ credible intervals from the hierarchical Bayesian analysis. Two binning schemes are presented: a coarse $5\times 5$ grid with $\Delta m=10\,M_{\odot}$ and $\Delta z=0.2$, and a finer $10\times 10$ grid with $\Delta m=5\,M_{\odot}$ and $\Delta z=0.1$. The latter exhibits larger uncertainties consistent with reduced event counts per bin. The linear term is consistent with zero for low mass systems ($m\lesssim 30\,M_{\odot}$) but increases toward higher masses, indicating massive black holes are relatively more abundant at earlier cosmic times. The quadratic term remains consistent with zero across all mass bins, implying a simple linear dependence on redshift adequately describes the current observational range $z\lesssim 1$ accessible to LVK.
To distinguish between genuine evolution of the BBH mass distribution and observational biases, we employ the LVK injection campaigns Observing Runs O1-O4. These datasets contain simulated signals with known source parameters, analyzed through the same search pipelines as the real data, thereby providing an empirical estimate of detection efficiency across the $(m,z)$ plane. For consistency with the observed catalog, we adopt a conservative network signal to noise threshold of $\rho_{\rm thresh}>10$ when defining detections in both the injections and the real sample. The resulting selection function, together with the full posterior samples of the observed events, is incorporated into our hierarchical Bayesian framework (Section 2). This framework models the intrinsic joint distribution $p(m,z)$ via a non parametric Taylor expansion, while explicitly marginalizing over measurement uncertainties and correcting for selection effects. The results of this analysis are presented in Section 4.
It is important to emphasize that Fig. 2 is exploratory only. The plotted medians and means provide intuition about the detected sample but do not constitute a statistically rigorous analysis, as they neglect selection effects and the full posterior structure of individual events.
4 Results
We apply the non parametric Bayesian framework described in Section 2 to the GWTC-3 and GWTC-4.0 catalog of BBH mergers, incorporating the corresponding injection campaigns to account for detector selection effects. The analysis is performed in the $(m,z)$ plane, discretized into grids of varying resolutions to study the impact of bin size on the results. We consider two cases: one with a $5\times 5$ grid characterized by $\Delta m=10\,M_{\odot}$ and $\Delta z=0.2$, and another with a finer $10\times 10$ grid having $\Delta m=5\,M_{\odot}$ and $\Delta z=0.1$. Both cases cover the redshift range $0\leq z\leq 1$ and mass range $30\,M_{\odot}\leq m\leq 80\,M_{\odot}$.
To ensure robustness, we restrict the analysis to systems with component masses above $30\,M_{\odot}$. At lower masses the GWTC-3 and GWTC-4.0 sample and the injection sets exhibit rapidly decreasing detection efficiency, sparse injection coverage, and larger relative mass uncertainties, which together render the inverse selection correction unstable. Limiting to $m\gtrsim 30\,M_{\odot}$ ensures that the inferred Taylor coefficients are driven by regions of the $(m,z)$ plane where both the observed events and the injections provide reliable information.
The detector selection function $S_{(}m,z)$ is calibrated from the injection campaign of LVK collaboration. For each bin $(m_{k},z_{j})$, we compute the detection probability as the fraction of injected signals recovered above the network SNR threshold (see Sec. 2), and then apply volume–time corrections to account for cosmological effects. The resulting selection function exhibits the expected strong dependence on both mass and redshift (Gerosa & Bellotti, 2024; Essick et al., 2025).
Figure 3 summarizes the key results of our non-parametric inference. The panels show the marginalized posterior distributions of the linear ($p_{1}$) and quadratic ($p_{2}$) evolution coefficients as functions of component mass. A clear trend emerges in the linear term: for lower mass systems ($m\lesssim 30\,M_{\odot}$), both $p_{1}(m)$ and $p_{2}(m)$ are consistent with zero, indicating no detectable redshift evolution. At higher masses, however, $p_{1}(m)$ shifts systematically toward positive values, suggesting that more massive black holes were relatively more common at earlier cosmic times. This pattern aligns with astrophysical models where metallicity influences the efficiency of forming massive remnants. Conversely, the quadratic coefficient $p_{2}(m)$ shows no statistically significant deviation from zero across all mass bins, implying that a simple linear and quadratic dependence on redshift sufficiently describes the currently accessible range ($z\lesssim 1$) probed by current LVK observations. It is important to also note that due to large number of bins present in this non-parametric model independent approach, the statistical uncertainties on the reconstruction is shadowing to draw any conclusive measurement of the redshift dependent mass distribution. A well motivated theoretical study with fewer number of parameters can unveil any potential trend with a better statistical significance.
These constraints are presented for two discretizations schemes in the $(m,z)$ plane: a coarser $5\times 5$ grid with $\Delta m=10\,M_{\odot}$, $\Delta z=0.2$ and a finer $10\times 10$ grid with $\Delta m=5\,M_{\odot}$, $\Delta z=0.1$. The finer binning reveals more detailed features in the evolution coefficients but comes with larger statistical uncertainties, as reflected in the increased error bars. This comparison highlights the trade-off between resolution and statistical significance when studying BBH mass evolution over redshift.
Taken together, the figure demonstrates that the observed BBH population is consistent with a scenario in which the abundance of massive black holes increases modestly with redshift, while lower mass systems remain approximately constant. The mass dependent nature of this evolution strongly points to environmental effects such as metallicity shaping the efficiency of massive black hole formation across cosmic time.
5 Conclusions
We have developed and applied a fully non-parametric Bayesian framework to infer the redshift evolution of the BBH mass distribution using all confident BBH detections from the LVK GWTC-3 and GWTC-4.0 catalogs.
By expanding the intrinsic joint distribution $p(m,z)$ as a Taylor series around $z_{\rm ref}=0.25$, the method captures evolutionary trends without imposing restrictive parametric assumptions, while rigorously accounting for detector selection effects through injection calibrated efficiencies.
Our analysis yields tentative but consistent evidence for mass dependent evolution. The linear evolution coefficient $p_{1}(m)$ is consistent with zero for lower mass systems ($m\lesssim 30\,M_{\odot}$), but becomes increasingly positive at higher masses. In particular, the highest mass bins show tentative evidence that massive black holes were relatively more abundant at earlier cosmic times. By contrast, the quadratic term $p_{2}(m)$ remains consistent with zero across all mass ranges, indicating that a linear description suffices up to the redshift range $z\sim 1$ currently probed by LVK observations.
These results are qualitatively consistent with theoretical expectations for metallicity driven evolution: low metallicity environments at high redshift suppress stellar winds, allowing the formation of more massive stellar cores and correspondingly heavier black hole remnants. The lack of detectable evolution for lower mass systems suggests that their formation channels are less sensitive to environmental variations, consistent with robust core collapse pathways. The approximately linear trend indicates gradual, metallicity driven changes over cosmic time rather than abrupt evolutionary transitions. This result agrees with the probability of different formation channels, we found in GWTC-4.0 using phase-space, showing that above approximately $45$ M there is an indication of a different population of BBHs (Afroz & Mukherjee, 2025c).
Beyond these specific findings, this work demonstrates the power of non parametric approaches for GW population inference. Unlike parametric models, which may obscure or bias subtle features, our Taylor expansion framework provides a flexible yet interpretable means of capturing population trends while fully correcting for observational selection effects (see Section 2). The method’s ability to reveal tentative mass dependent evolution highlights its value for future analyses with larger datasets.
Previous studies based on GWTC-3 (Fishbach et al., 2021; van Son et al., 2022; Tiwari, 2022; Karathanasis et al., 2023; Rinaldi et al., 2024; Heinzel et al., 2025b; Lalleman et al., 2025) reported mixed evidence for mass–redshift evolution, ranging from no detectable trends to weak indications at high masses. By incorporating the larger GWTC-4.0 along with GWTC-3 catalog, our analysis reveals tentative evidence for mass-dependent linear evolution.
Looking forward, the dramatic increase in detection rates expected from next generation ground based detectors such as Cosmic Explorer (Reitze et al., 2019) and the Einstein Telescope (Branchesi et al., 2023) will extend BBH catalogs to higher redshift, enabling reconstruction of higher order evolutionary features with high statistical precision. Our non parametric methodology is well positioned to exploit these datasets, offering both robustness and flexibility in probing the cosmic history of black hole formation. As GW astronomy transitions into the precision era, such model independent frameworks will be essential for extracting the full astrophysical information encoded in BBH populations, ultimately linking GW observations to stellar evolution, galactic chemical enrichment, and the early universe.
Data Availability
The gravitational wave catalogs used in this study are publicly available on Zenodo as part of the LVK data releases: GWTC-2.1 (link), GWTC-3 (link) and GWTC-4.0 (link).
Acknowledgments
The authors express their gratitude to Souradeep Pal for reviewing the manuscript and providing useful comments as a part of the LIGO publication policy. This work is part of the ⟨data|theory⟩ Universe Lab, supported by TIFR and the Department of Atomic Energy, Government of India. The authors express gratitude to the system administrator of the computer cluster of ⟨data|theory⟩ Universe Lab. This research is supported by the Prime Minister Early Career Research Award, Anusandhan National Research Foundation, Government of India. We thank the LIGO Scientific Collaboration, the Virgo Collaboration, and the KAGRA Collaboration for providing the gravitational-wave data through the Gravitational Wave Open Science Center (GWOSC). LIGO, funded by the U.S. National Science Foundation (NSF), and Virgo, supported by the French CNRS, Italian INFN, and Dutch Nikhef, along with contributions from Polish and Hungarian institutes. This collaborative effort is backed by the NSF’s LIGO Laboratory, a major facility fully funded by the National Science Foundation. The research leverages data and software from the Gravitational Wave Open Science Center, a service provided by LIGO Laboratory, the LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA. Advanced LIGO’s construction and operation receive support from STFC of the UK, Max Planck Society (MPS), and the State of Niedersachsen/Germany, with additional backing from the Australian Research Council. Virgo, affiliated with the European Gravitational Observatory (EGO), secures funding through contributions from various European institutions. Meanwhile, KAGRA’s construction and operation are funded by MEXT, JSPS, NRF, MSIT, AS, and MoST. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge the use of the following packages in this work: Numpy (Van Der Walt et al., 2011), Scipy (Jones et al., 2001–), Matplotlib (Hunter, 2007), Astropy (Robitaille et al., 2013) and emcee (Foreman-Mackey et al., 2013).